OpenStudy (anonymous):

Use the associative property to Identify which expression is equal to (54)(12)(2x) A. (54)(12)(2x) B. (54*12*2)(x) C. 54(12+2x) D. (2x)(12)(54)

OpenStudy (anonymous):

I need help ): im in corse recovery and i cant fail

OpenStudy (anonymous):

will give medal (:

OpenStudy (anonymous):

Google -> Associative property

OpenStudy (anonymous):

i did and its doesnt have any with the parenthesis like that

OpenStudy (anonymous):

I dunno, I typed that into Google and the very first link was to wikipedia which contained plenty of examples.

OpenStudy (anonymous):

well did you get the answer from the examples?

OpenStudy (anonymous):

No, I know the answer because I know what associativity means. But from the examples in the second link from the search you can certainly work out the answer.

OpenStudy (anonymous):

Well i looked and im not sure where to find it and this is a timed test ):

OpenStudy (anonymous):

This is the second link http://math.about.com/od/prealgebra/a/associative.htm

OpenStudy (anonymous):

so would the answer be D? @Bad2zBone

OpenStudy (anonymous):

I will explain this and people will probably do that here for you but you should learn to research information by yourself. You will NEVER learn by using this website for answers. Multiplication and addition are associative. That means you can group objects any way you like and this will not change the outcome. By grouping i mean setting bracekts '()'. (3 x 1) x 5 = 3 x (1 x 5) or 3 + (1 + 4) + 3 = 3 + 1 + (4 + 3) Addition and multiplication is also COMMUTATIVE. Which means that the order in which you do the operation of multiplication or addition does not matter. So: 3 + 1 + 5 = 1 + 5 + 3 1 x 5 x 3 = 5 x 1 x 1

OpenStudy (anonymous):

That last one should be 1 x 5 x 3 = 5 x 1 x 3

OpenStudy (anonymous):

Okay well i appreciate the help but im still curious if i got the answer correct or not, i dont want to turn in the test and it be wrong when it is also time'd.

OpenStudy (anonymous):

You have 54 * 12 * 2x. By rearranging this into 2x * 12 * 54 which property are you using?

OpenStudy (anonymous):

Associative property if im not mistaken

OpenStudy (anonymous):

So you haven't read anything i've posted above then?

OpenStudy (anonymous):

I Read all of it but frankly i am confused and it would be communative but that still doesnt help me out with the answer

OpenStudy (anonymous):

You have 54 * 12 * 2x. By rearranging this into 2x * 12 * 54 you are using the commutative property since you've swapped the order of operations. 54(12+2x) This is clearly incorrect. A. (54)(12)(2x) Nothing is used here. So you had 54 * 12 * 2x, you insert brackets (group stuff together) to get this: (54*12*2)(x). By grouping 54,12,2 in one group and x in the other group you have used the associative property

OpenStudy (anonymous):

Okay but you just gave me two answer's of the associative property and those are both on the test im geussing 54(12+2x) is the correct one though even though you said both were correct

OpenStudy (anonymous):

@Bad2zBone

OpenStudy (anonymous):

54(12+2x) this is not equal to (54)(12)(2x) so it is incorrect.

OpenStudy (anonymous):

okay i understand now